OpenStudy (mtalhahassan2):

Need Help! Convert each angle from radians to degrees. Express the measure correct to two decimal places, if necessary. a) 3

Nnesha (nnesha):

\[\huge\rm r \times \frac{ \pi }{ 180 }\] multiply Radian by pi over 180 remember \[\rm \pi = 180 \] degree

OpenStudy (mtalhahassan2):

so what would be the answer be

Nnesha (nnesha):

r = ?? pi = 3.14 or just put that into the calculator and solve let me know what you get!

OpenStudy (mtalhahassan2):

0.0174444444

OpenStudy (mtalhahassan2):

what you get

OpenStudy (mtalhahassan2):

3.14/180

Nnesha (nnesha):

r = radian which is 3 in the question

OpenStudy (mtalhahassan2):

0.52359877

Nnesha (nnesha):

wait a sec......hm

OpenStudy (mtalhahassan2):

3(3.14) divide 180

Nnesha (nnesha):

is that just a 3 or 3 degree ?

OpenStudy (mtalhahassan2):

just 3

Nnesha (nnesha):

ohh okay so that just a 3 that's why first you have to change that into pi by dividing 3/pi

Nnesha (nnesha):

and then convert to degree

OpenStudy (mtalhahassan2):

oh ok

OpenStudy (mtalhahassan2):

171.97

Nnesha (nnesha):

did you divide by pi

OpenStudy (butterflydreamer):

hmm.. when we convert radians to degrees we do: